Suppose the government passe a law to keep price above the $16 .

*How many units of this good will be bought and sold .?

*Calculate the value to consumers at this quantity?

* Calculate the consumers surplus at this quantity ?

* Calculate the cost of producing this quantity ?

* Calculate the producer surplus at this quantity ?

* Calculate the change in consumer surplus caused by this price floor ?

*Calculate the changenin producer surplus caused by this price floor?

* Calculate the dead weight loss caused by this price floor ?


1
Expert's answer
2022-01-19T16:16:51-0500

Admit

Qd=P-3

Qs=30-P

P=17

Qd=17-3=14

QS=30-17=13


Qd=Qs

P-3=30-P

2P=33

P=16.5

Q=13.5


"16.5\times13.5=222.75"


"consumers surplus=0.5Qd\Delta P=0.5\times13.5(17-16.5)=3.375"

Pmin=15

"producer surplus=0.5\times(16.5-15)\times13.5=10.125"

"change in consumer surplus caused=-3.375-0.5\times13.5(17-15)=-10,125"

"change in producer surplus=10.125-0.5\times13.5(17-15)=23.625"

"dead weight loss=0.5\times*(17-15)(13.5-13)=0.5"